# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 2.3. Orbifold Chern classes
where the first equality follows from the definition 2.13 and Lemma 2.17.
Let $\mathcal{Q}_{\mathrm{orb}}:=\mathcal{E}_{\mathrm{orb}}/\mathcal{F}_{\mathrm{orb}}$. $\mathcal{T}_{\mathrm{orb}}:=(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}/\mathcal{Q}_{\mathrm{orb}})$ is a torsion orbi-sheaf with the support $S_{\mathrm{orb}}$. For any open subset $V$ compactly supported in $Y\setminus g^{-1}(Z)$, Proposition 2.19 implies that there exists a smooth orbifold semi-positive $(1,1)$-form $\omega_{i,\mathrm{orb}}\in[g_{\mathrm{orb}}^{*}\omega_{i}]$ such that $\omega_{i,\mathrm{orb}}$ is strictly positive on $V_{\mathrm{orb}}$. Building on the GRR formula for orbifold embeddings (cf. [78, Theorem 1.1]) and the fact that $\mathrm{codim}(S_{\text{orb}})\geq 2$, it follows that
(2.4) $\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}/\mathcal{Q}_{\mathrm{orb}})\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]\geq 0.$
@@
Moreover, equality forces $V_{\text{orb}}\cap S_{\text{orb}}=\emptyset$, as can be shown by adapting the reasoning in [5, Section 7]. This is shown as follows. First, we take an orbifold embedded resolution $\pi_{\mathrm{orb}}:(\widehat{X}_{\mathrm{orb}},\widehat{\mathcal{S}}_{\mathrm{orb}})\rightarrow(X_{\text{orb}},Z_{\text{orb}})$ of $(X_{\text{orb}},Z_{\text{orb}})$ and all irreducible components $\widehat{\mathcal{S}}_{i,\text{orb}}$ of $\mathcal{S}_{\text{orb}}$ smooth with transversal intersection, which exists by the functorial embedded resolution [105, Theorem 2.0.2]. Then, an application of the GRR formula for orbifold embeddings gives
@@
$\pi^{*}\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{T}_{\mathrm{orb}})\cdot\pi^{*}[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots\pi^{*}[g_{\mathrm{orb}}^{*}\omega_{n-2}]=\sum\limits\operatorname{rank}(\mathcal{T}_{\mathrm{orb}}|_{{\mathcal{S}}_{i,\mathrm{orb}}})\int_{\mathcal{S}_{i,\mathrm{orb}}}\pi_{\mathrm{orb}}^{*}\omega_{1,\mathrm{orb}}\wedge\cdots\wedge\omega_{n-2,\mathrm{orb}}$
@@
since the image of the intersection of $\widehat{\mathcal{S}}_{\mathrm{orb}}$ and the exceptional divisor of $\pi_{\mathrm{orb}}$ has codimension at least $3$.
Combining (2.3), (2.4), we obtain
@@
$(\widehat{\mathrm{ch}}_{2}(\mathcal{E})-\widehat{\mathrm{ch}}_{2}(\mathcal{F}))\cdot[\omega_{1}]\cdots[\omega_{n-2}]=$ $(-\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}/\mathcal{Q}_{\mathrm{orb}})+\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}))\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]$
$\leq$ $\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee})\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]$
$=$ $\widehat{\mathrm{ch}}_{2}(\mathcal{Q}^{\vee\vee})\cdot[\omega_{1}]\cdots[\omega_{n-2}]$
@@
where the last equality follows from the fact that $\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}=g_{\mathrm{orb}}^{\vee\vee}\mathcal{Q}^{\vee\vee}$ outside $g_{\mathrm{orb}}^{-1}(Z)$. Then the inequality was proved. Moreover, equality forces $(Y\setminus f^{-1}(Z))_{\mathrm{orb}}\cap S_{\mathrm{orb}}=\emptyset$ since the choice of $V$ is arbitrary. This implies that $0\rightarrow\mathcal{F}_{\mathrm{orb}}\rightarrow\mathcal{E}_{\mathrm{orb}}\rightarrow\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}\rightarrow 0$ is a short exact sequence of orbi-sheaves over $Y\setminus f^{-1}(Z)$, and hence it induces a short exact of reflexive sheaves
$0\rightarrow\mathcal{F}\rightarrow\mathcal{E}\rightarrow\mathcal{Q}^{\vee\vee}\rightarrow 0$
on $X\setminus Z$ by [33, Lemma A.3 and Lemma A.4]. In particular, $\mathcal{Q}^{\vee\vee}=\mathcal{Q}$ outside $Z$. The proof is complete. ∎
@@ 2.4. Quasi-étale morphisms
Let $X$ be a compact klt space of dimension $n$ and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ and $\gamma:X^{\prime}\rightarrow X$ is a quasi-étale Galois morphism between complex spaces induced by a finite group $G\subset\mathrm{Aut}(X^{\prime})$. Then $Y$ is also klt and the following equality holds for any reflexive sheaf $\mathcal{E}$ and any classes $\sigma\in H_{2n-2}(X,\mathbb{R})$
$\widehat{c}_{2}(\gamma^{[*]}\mathcal{E})\cdot\gamma^{*}\sigma=\deg\gamma\cdot\widehat{c}_{2}(\mathcal{E})\cdot\sigma$
Ditto for $\widehat{c}_{1}(\mathcal{E})$ and $\widehat{c}_{1}^{2}(\mathcal{E})$.
@@
The fact that $X^{\prime}$ is klt follows from the argument of [66, Proposition 5.20]. Let $f:Y\rightarrow X$, $f^{\prime}:Y^{\prime}\rightarrow X^{\prime}$ be partial orbifold resolutions of $X,X^{\prime}$, respectively and let $Y_{\mathrm{orb}}:=\{V_{\alpha}^{\prime},G_{\alpha}^{\prime},\mu_{\alpha}^{\prime}\},Y_{\mathrm{orb}}^{\prime}:=\{V_{\beta}^{\prime},G_{\beta}^{\prime},\mu_{\beta}^{\prime}\}$ be the maximal standard orbifold structures of $Y^{\prime},Y$, respectively. Denote the adjoint of the indeternimacy locus of $f^{-1}$ and the image of the indeterminacy locus of $(f^{\prime})^{-1}$ under $\gamma$ by $Z$, which has codimension at least $3$. Recall Lemma 2.18, we may assume that there exists some orbi-bundles $E_{\mathrm{orb}},E_{\mathrm{orb}}^{\prime}$ on $Y_{\mathrm{orb}},Y_{\mathrm{orb}}^{\prime}$ such that $E_{\mathrm{orb}}=f_{\mathrm{orb}}^{*}\mathcal{E},E_{\mathrm{orb}}^{\prime}=f_{\mathrm{orb}}^{\prime*}\mathcal{E}^{\prime}$ outside $f^{-1}(Z),(\gamma\circ f^{\prime})^{-1}(Z)$, respectively, and
@@
$c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot f^{*}\sigma=\deg(\gamma)\cdot c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}}^{\prime})\cdot(f^{\prime})^{*}\gamma^{*}\sigma.$
Let $g_{\alpha,\beta}=W_{\alpha,\beta}\rightarrow V_{\alpha}\times_{X}V_{\beta}^{\prime}$ be a functorial resolution of singularities (cf. [22, Theorem 3.10]), then there is a induced action $G_{\alpha}\times G_{\beta}$ on $W_{\alpha,\beta}$ such that $g_{\alpha,\beta}$ is $(G_{\alpha}\times G_{\beta})$-invariant, which induces an effective complex orbifold $W_{\mathrm{orb}}:=\{W_{\alpha,\beta},G_{\alpha}\times G_{\beta}\}$. Let $W$ be the underlying space and we have the following commute diagram
@@
Since $\gamma$ is quasi-étale, $\{V_{\beta}^{\prime},GG_{\beta}^{\prime},h\circ\mu_{\beta}^{\prime}\}$ can be viewed as a non-maximal standard orbifold structure of $Y\setminus\gamma^{-1}(Z)$. Applying Lemma 2.1, we conclude that the singularities of $Y^{\prime}\times_{X}Y$ lies in $pr_{2}^{-1}(\gamma^{-1}(Z))$, and thus Lemma 2.17 implies that
$c_{2}^{\mathrm{orb}}(g^{*}pr_{2}^{*}E_{\mathrm{orb}})\cdot g^{*}pr_{2}^{*}f^{*}\sigma=c_{2}^{\mathrm{orb}}(g^{*})\cdot c_{2}^{\mathrm{orb}}(g^{*}pr_{1}^{*}E_{\mathrm{orb}}^{\prime})\cdot g^{*}pr_{1}^{*}f^{*}\sigma$
because $g^{*}pr_{2}^{*}E_{\mathrm{orb}}$ and $g^{*}pr_{2}^{*}E_{\mathrm{orb}}^{\prime}$ coincides outside the preimage of $Z$. Note that $pr_{1}$ is generically biholomorphic and $pr_{2}$ is finite with degree $\deg(\gamma)$. The proof is complete. ∎
@@
Using a similar argument of the proof of Proposition 2.21, we have that for any locally free sheaf $\mathcal{E}$, $\widehat{c}_{1}(\mathcal{E}),\widehat{c}_{1}^{2}(\mathcal{E}),\widehat{c}_{2}(\mathcal{E})$ equal the usual product with the topological Chern classes of $\mathcal{E}$ in the sense of homology class.
@@
Let $\mathcal{E}$ be a torsion-free sheaf on a compact normal space $X$. Then the following statements holds.
@@
    $c_{1}(\gamma^{*}\mathcal{F})\cdot\gamma^{*}\sigma=\deg\gamma\cdot c_{1}(\mathcal{F})\cdot\gamma.$
@@
The statement (1) can be implied by a similar argument of Lemma 2.21. The statement (2) follows from that for any partial orbifold resolution of singularities $f:Y\rightarrow X$, $f_{*}c_{1}(f^{[*]}\mathcal{F})=c_{1}(\mathcal{F})$ (see e.g. [110, Lemma 2.31]) and Lemma [110, Lemma 2.37]. ∎
@@
    $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\alpha_{0},\cdots,\alpha_{n-2})$-stable if and only if $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $G$-stable with respect to $\gamma^{*}\alpha_{0},\cdots,\gamma^{*}\alpha_{n-2}$. Ditto for polystability.
@@
Based on the above lemmas, the statement (1) follows directly from the similar argument [110, Proposition 4.14]. Regarding (2), it suffices to notice that the uniqueness of the maximal destablizing subsheaf $\mathcal{F}_{Y_{\mathrm{reg}}}$ of $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ implies that $\mathcal{F}_{Y_{\mathrm{reg}}}$ is $G$-equivariant and any $G$-equivariant $\gamma^{[*]}\theta_{X_{\mathrm{reg}}}$-invariant subsheaf of $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ descends to a $\theta_{X_{\mathrm{reg}}}$ subsheaf of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$. ∎
3. Pluri-harmonic metrics of Higgs sheaves over the regular loci

### v2

@@ 2.3. Orbifold Chern classes
where the first equality follows from Definition 2.13 and Lemma 2.17.
Let $\mathcal{Q}_{\mathrm{orb}}:=\mathcal{E}_{\mathrm{orb}}/\mathcal{F}_{\mathrm{orb}}$. Then $\mathcal{T}_{\mathrm{orb}}:=\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}/\mathcal{Q}_{\mathrm{orb}}$ is a torsion orbi-sheaf supported in some orbifold analytic subspace $S_{\mathrm{orb}}$. Based on the GRR formula for orbifold embeddings (cf. [79, Theorem 1.1]) and the fact that $\mathrm{codim}(S_{\text{orb}})\geq 2$, it follows that
(2.4) $\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{T}_{\mathrm{orb}})\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]\geq 0,$
@@
as can be shown by adapting the reasoning in [5, Section 7]. This is shown as follows. First, we take an orbifold embedded resolution $\pi_{\mathrm{orb}}:(\widehat{W}_{\mathrm{orb}},\widehat{S}_{\mathrm{orb}})\rightarrow(W_{\text{orb}},S_{\text{orb}})$ such that all irreducible components $\widehat{S}_{i,\text{orb}}$ of $\widehat{S}_{\text{orb}}$ smooth with transversal intersection, which exists by the functorial embedded resolution (cf. [108, Theorem 2.0.2]). Then, an application of the GRR formula for orbifold embeddings gives
@@
(2.5) $\begin{split}&\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{T}_{\mathrm{orb}})\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]=\pi^{*}\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{T}_{\mathrm{orb}})\cdot\pi^{*}[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots\pi^{*}[g_{\mathrm{orb}}^{*}\omega_{n-2}]\\ &=\sum\limits\operatorname{rank}(\mathcal{T}_{\mathrm{orb}}|_{{\widehat{S}}_{i,\mathrm{orb}}})\int_{{\widehat{S}}_{i,\mathrm{orb}}}\pi_{\mathrm{orb}}^{*}g_{\mathrm{orb}}^{*}\omega_{1}\wedge\cdots\wedge\pi_{\mathrm{orb}}^{*}g_{\mathrm{orb}}^{*}\omega_{n-2}\\ &=\sum\limits\operatorname{rank}(\mathcal{T}_{\mathrm{orb}}|_{{\widehat{S}}_{i,\mathrm{orb}}})\int_{{S}_{i,\mathrm{orb}}}g_{\mathrm{orb}}^{*}\omega_{1}\wedge\cdots\wedge g_{\mathrm{orb}}^{*}\omega_{n-2}\geq 0\end{split}$
@@
since the image of the intersection of $\widehat{S}_{\mathrm{orb}}$ and the exceptional divisor of $\pi_{\mathrm{orb}}$ has codimension at least $3$, where $\pi:\widehat{W}\rightarrow W$ is the holomorphic map between the underlying complex spaces induced by $\pi_{\mathrm{orb}}$.
Combining (2.3) and (2.5), we obtain
@@
(2.6) $\begin{split}(\widehat{\mathrm{ch}}_{2}(\mathcal{E})-\widehat{\mathrm{ch}}_{2}(\mathcal{F}))\cdot[\omega_{1}]\cdots[\omega_{n-2}]=&(-\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{T}_{\mathrm{orb}})+\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}))\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]\\ \leq&\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee})\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]\\ =&\widehat{\mathrm{ch}}_{2}(\mathcal{Q}^{\vee\vee})\cdot[\omega_{1}]\cdots[\omega_{n-2}],\end{split}$
@@
where the last equality follows from the fact that $\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}=g_{\mathrm{orb}}^{[*]}\mathcal{Q}^{\vee\vee}$ outside $g_{\mathrm{orb}}^{-1}(Z)$. Then the inequality (2.2) was proved.
Now we assume that $\omega_{0},\cdots,\omega_{n-2}$ are Kähler forms and that the equality in (2.2) holds. Then for any open subset $V$ compactly contained in $W\setminus g^{-1}(Z)$, Proposition 2.19 means that there exists smooth orbifold semipositive $(1,1)$-forms $\omega_{1,\mathrm{orb}}\in[g_{\mathrm{orb}}^{*}\omega_{1}],\cdots,\omega_{n-2,\mathrm{orb}}\in[g_{\mathrm{orb}}^{*}\omega_{n-2}]$ on $W_{\mathrm{orb}}$ which are strictly positive on $g_{\mathrm{orb}}^{-1}(V)$. Moreover, the above argument implies that the inequality in the third line of (2.5) is in fact an equality, which yields that
$\omega_{1,\mathrm{orb}}\wedge\cdots\wedge\omega_{n-2,\mathrm{orb}}=0$
on each $S_{i,\mathrm{orb}}$. Hence $S_{i,\mathrm{orb}}\cap g_{\mathrm{orb}}^{-1}(V)=\emptyset$. Since $V$ is arbitrary, we conclude that $\left(W_{\mathrm{orb}}\setminus g_{\mathrm{orb}}^{-1}(Z)\right)\cap S_{\mathrm{orb}}=\emptyset$ and consequently $\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}=\mathcal{Q}_{\mathrm{orb}}$ on $W_{\mathrm{orb}}\setminus g_{\mathrm{orb}}^{-1}(Z)$. This gives rise to a short exact sequence of orbi-sheaves $0\rightarrow\mathcal{F}_{\mathrm{orb}}\rightarrow\mathcal{E}_{\mathrm{orb}}\rightarrow\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}\rightarrow 0$ over $W_{\mathrm{orb}}\setminus g_{\mathrm{orb}}^{-1}(Z)$. By [33, Lemma A.3 and Lemma A.4], it induces a short exact sequence of reflexive sheaves $0\rightarrow\mathcal{F}\rightarrow\mathcal{E}\rightarrow\mathcal{Q}^{\vee\vee}\rightarrow 0$ on $X\setminus Z$. In particular, $\mathcal{Q}^{\vee\vee}=\mathcal{Q}$ outside $Z$. This completes the proof. ∎
@@ 2.4. Quasi-étale morphisms
Let $X$ be a compact klt space of dimension $n$ and $\gamma:X^{\prime}\rightarrow X$ be a quasi-étale Galois morphism induced by a finite group $G\subset\mathrm{Aut}(X^{\prime})$. Then $X^{\prime}$ is also klt, and for any reflexive sheaf $\mathcal{E}$ on $X$ and any class $\sigma\in H_{2n-2}(X,\mathbb{R})$, we have
$\widehat{c}_{2}(\gamma^{[*]}\mathcal{E})\cdot\gamma^{*}\sigma=\deg\gamma\cdot\widehat{c}_{2}(\mathcal{E})\cdot\sigma.$
The same holds for $\widehat{c}_{1}(\mathcal{E})$ and $\widehat{c}_{1}^{2}(\mathcal{E})$.
@@
The fact that $X^{\prime}$ is klt follows from the argument of [67, Proposition 5.20]. Suppose $f:Y\rightarrow X$ and $f^{\prime}:Y^{\prime}\rightarrow X^{\prime}$ are partial orbifold resolutions of $X$ and $X^{\prime}$, respectively. Let $Y_{\mathrm{orb}}:=\{(V_{\alpha},G_{\alpha},\mu_{\alpha})\}$ and $Y_{\mathrm{orb}}^{\prime}:=\{(V_{\beta}^{\prime},G_{\beta}^{\prime},\mu_{\beta}^{\prime})\}$ be the maximal standard orbifold structures of $Y$ and $Y^{\prime}$, respectively. Denote by $Z$ the union of the indeterminacy locus of $f^{-1}$ and the image under $\gamma$ of the indeterminacy locus of $(f^{\prime})^{-1}$, a set which has codimension at least $3$. Recalling Lemma 2.18, we may assume that there exist orbi-bundles $E_{\mathrm{orb}}$ on $Y_{\mathrm{orb}}$ and $E_{\mathrm{orb}}^{\prime}$ on $Y_{\mathrm{orb}}^{\prime}$ such that $E_{\mathrm{orb}}=f_{\mathrm{orb}}^{*}\mathcal{E}$ on $Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(Z)$ and $E_{\mathrm{orb}}^{\prime}=(f_{\mathrm{orb}}^{\prime})^{*}\gamma^{[*]}\mathcal{E}$ on $Y_{\mathrm{orb}}^{\prime}\setminus{(\gamma\circ f^{\prime})}_{\mathrm{orb}}^{-1}(Z)$, and that the following equalities hold:
@@
$\deg(\gamma)\cdot c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot f^{*}\sigma=c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}}^{\prime})\cdot(f^{\prime})^{*}\gamma^{*}\sigma.$
Let $g_{\alpha,\beta}:=W_{\alpha,\beta}\rightarrow V_{\alpha}\times_{X}V_{\beta}^{\prime}$ be a functorial resolution of singularities (cf. [22, Theorem 3.10]), then there is an induced action $G_{\alpha}\times G_{\beta}^{\prime}$ on $W_{\alpha,\beta}$ such that $g_{\alpha,\beta}$ is $(G_{\alpha}\times G_{\beta}^{\prime})$-invariant, which induces an effective complex orbifold $W_{\mathrm{orb}}:=\{(W_{\alpha,\beta},G_{\alpha}\times G_{\beta}^{\prime})\}$. Let $W$ be the underlying space and we have the following commutative diagram
@@
where $g$ is the morphism induced by $g_{\alpha,\beta}$. Since $\gamma$ is quasi-étale, $\{(V_{\beta}^{\prime},GG_{\beta}^{\prime},f^{-1}\circ\gamma\circ f^{\prime}\circ\mu_{\beta}^{\prime})\}$ is a non-maximal standard orbifold structure of $Y\setminus f^{-1}(Z)$ with the natural group action $GG_{\beta}^{\prime}$. Applying Lemma 2.1, we conclude that $V_{\alpha}^{\prime}\times_{X}V_{\beta}^{\prime}$ is smooth over $\left(Y\times_{X}Y^{\prime}\right)\setminus\left(pr_{1}^{-1}(f^{-1}(Z))\right)$ and thus the indeterminacy locus of $g^{-1}$ is contained in $pr_{1}^{-1}(f^{-1}(Z))$. Let $g_{\mathrm{orb},1}:W_{\mathrm{orb}}\rightarrow Y_{\mathrm{orb}}$ and $g_{\mathrm{orb},2}:W_{\mathrm{orb}}\rightarrow Y_{\mathrm{orb}}^{\prime}$ be the orbifold holomorphisms induced by $pr_{1}\circ g$ and $pr_{2}\circ g$, respectively. Therefore, Lemma 2.17 implies that
$c_{2}^{\mathrm{orb}}(g_{\mathrm{orb},2}^{*}E_{\mathrm{orb}}^{\prime})\cdot g^{*}pr_{2}^{*}(f^{\prime})^{*}\gamma^{*}\sigma=c_{2}^{\mathrm{orb}}(g_{\mathrm{orb},1}^{*}E_{\mathrm{orb}})\cdot g^{*}pr_{1}^{*}f^{*}\sigma$
because $g_{\mathrm{orb},1}^{*}E_{\mathrm{orb}}$ and $g_{\mathrm{orb},2}^{*}E_{\mathrm{orb}}^{\prime}$ coincide outside the preimage of $Z$. Note that $pr_{2}\circ g$ is generically biholomorphic, and $pr_{1}\circ g$ is generically finite and étale with degree $\deg(\gamma)$. The proof is complete. ∎
@@
Using an argument similar to that in the proof of Proposition 2.21, we find that for any locally free sheaf $\mathcal{E}$, the classes $\widehat{c}_{1}(\mathcal{E}),\widehat{c}_{1}^{2}(\mathcal{E})$ and $\widehat{c}_{2}(\mathcal{E})$ coincide, respectively, with the topological Chern classes of $\mathcal{E}$ as elements of the homology groups.
@@
Let $\mathcal{F}$ be a torsion-free sheaf on a compact normal space $X$. Then the following statements hold.
@@
    $c_{1}(\gamma^{*}\mathcal{F})\cdot\gamma^{*}\sigma=\deg\gamma\cdot c_{1}(\mathcal{F})\cdot\sigma.$
@@
Statement (1) can be deduced by an argument similar to that of Proposition 2.21. Statement (2) follows from the fact that for any orbifold modification of singularities $f:Y\rightarrow X$, one has $f_{*}c_{1}(f^{[*]}\mathcal{F})=c_{1}(\mathcal{F})$ (see e.g. [113, Lemma 2.31]), together with [113, Lemma 2.37]. ∎
@@
    $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\alpha_{0},\cdots,\alpha_{n-2})$-stable if and only if $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $G$-stable with respect to $\gamma^{*}\alpha_{0},\cdots,\gamma^{*}\alpha_{n-2}$. The same holds for polystability.
@@
Based on the above lemmas, statement (1) follows directly from an argument similar to that of [113, Proposition 4.14]. Regarding (2), it suffices to notice that the uniqueness of the maximal destablizing subsheaf $\mathcal{F}_{Y_{\mathrm{reg}}}$ of $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ implies that $\mathcal{F}_{Y_{\mathrm{reg}}}$ is $G$-equivariant and any $G$-equivariant $\gamma^{[*]}\theta_{X_{\mathrm{reg}}}$-invariant subsheaf of $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ descends to a $\theta_{X_{\mathrm{reg}}}$-invariant subsheaf of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$. ∎
3. Pluri-harmonic metrics on Higgs sheaves over the regular locus